A geiger-muller tube is a type of gas-filled radiation detector. It can detect particles like X-rays, alpha particles and beta rays (electrons). This is useful in quantizing the activity of a radioactive source or determining if an area containing radioactive material is safe to enter. Assuming that you have one mole of as, if a geiger counter is filled with 9674 Pa of argon gas at room tempae (T ensity of the gas in this geiger tube? 25.5 C), what is the Number g/cm
Answer #1

PV = nRT

n = number of moles

R = gas constant = 0.08206 (atm∙L)/(mol∙K)

T = temperature in kelvins

P = pressure in atm (1 atm = 101 kPa)

V = volume in liters

P = 9.674 kPa = 9.674/101 = 0.09578 atm

T = 25.5+273 K = 298.5 K

V = nRT/P = (1)(0.08206)(298.5) / (0.09578) = 255.741 L or 255741 cm3

1 mole of argon = 40g

so density is 40g / 255741cm3 = 1.56*10^-4 g/cm^3